# CoAR: Concept Injection into Autoregressive Models for Personalized Text-to-Image Generation

### v1

@@ Abstract
The unified autoregressive (AR) model excels at multimodal understanding and generation, but its potential for customized image generation remains underexplored. Existing customized generation methods rely on full fine-tuning or adapters, making them costly and prone to overfitting or catastrophic forgetting. In this paper, we propose CoAR, a novel framework for injecting subject concepts into the unified AR models while keeping all pre-trained parameters completely frozen. CoAR learns effective, specific subject representations with only a minimal number of parameters using a Layerwise Multimodal Context Learning strategy. To address overfitting and language drift, we further introduce regularization that preserves the pre-trained distribution and anchors context tokens to improve subject fidelity and re-contextualization. Additionally, CoAR supports training-free subject customization in a user-provided style. Experiments demonstrate that CoAR achieves superior performance on both subject-driven personalization and style personalization, while delivering significant gains in computational and memory efficiency. Notably, CoAR tunes less than 0.05% of the parameters while achieving competitive performance compared to recent Proxy-Tuning. Code: https://github.com/KZF-kzf/CoAR.
@@
Figure 1: Visualization of samples generated by our proposed CoAR, showcasing subject personalization (left), where a few reference images are provided, and style personalization (right), where the subject is rendered in user-specified artistic styles without additional training.
Introduction
Subject-driven image generation (Chen et al. 2023b; Ruiz et al. 2023; Li, Li, and Hoi 2023; Shin et al. 2025) aims to generate images of a specific subject or concept in new scenes, given only a few reference images. This capability has found applications in virtual try-on (Ye et al. 2025; Zhang et al. 2024), personalized content creation (Liu et al. 2024b), and advertising design (Chen et al. 2023b).
Most of the existing approaches, such as DreamBooth (Ruiz et al. 2023) and its extensions (Chen et al. 2023a; Han et al. 2023; Kumari et al. 2023), customize generation by adapting the distribution of pre-trained diffusion models to match the subject distribution through full fine-tuning or parameter-efficient techniques like LoRA (Hu et al. 2022). These methods typically associate the subject with a unique identifier [V], enabling personalized generation via prompt injection during inference. However, such techniques are predominantly designed for diffusion models (Esser et al. 2024; Rombach et al. 2022; Podell et al. 2023) and are not directly transferable to multimodal autoregressive (AR) models. Applying these methods naively to AR models poses two challenges.
  • Risk of overfitting. AR models are pretrained on large-scale multimodal corpora and expected to support both understanding and generation. In contrast, the target subject typically has only a few reference images (3–5). Such scarce data makes it extremely difficult to adapt the pre-trained models to the subject distribution, resulting in a risk of overfitting and catastrophic forgetting of the pre-trained knowledge.
  • Inefficiency. Extensively adjusting the pretrained model parameters to learn a single concept with the limited number of images per subject is intuitively inefficient.
We argue that instead of fine-tuning the pretrained model on limited data to shift its distribution toward a target subject, it is more effective to inject subject concepts directly into the model’s existing distribution, thereby fully leveraging the generative capabilities of large models. To this end, we introduce Concept Injection for AR Models (CoAR), an extremely efficient framework for customized generation in multimodal AR models. CoAR adopts a Layer-wise Multimodal Context Learning (LMCL) strategy, encoding subject-specific information through a small number of learnable prompts without modifying any pretrained parameters. To guide learning while preserving the model’s original capabilities, we propose a Dual Prior Preservation (DPP) loss that constrains training. Additionally, we design a Context-Aware Self-Regularization (CASR) loss to improve re-contextualization stability and output quality. Built upon this foundation, CoAR also supports training-free customization, enabling arbitrary subject–style combinations during inference while maintaining strong re-contextualization performance. With a frozen backbone and only minimal parameter tuning, CoAR effectively injects the specific subject concept into the AR model and mitigates overfitting through multi-faceted regularization.
Extensive experiments demonstrate the superiority of CoAR in subject fidelity, instruction alignment, and parameter efficiency. Compared to naive DreamBooth adaptation in AR models, CoAR achieves significant performance gains while training approximately fewer than 0.1M parameters. Against existing AR-based customization methods, CoAR achieves competitive or superior results while training with no more than 0.15% of the parameters used by other approaches. For stylization tasks, our training-free method achieves subject identity preservation and high-fidelity stylization across diverse scenes, matching the modern LoRA-based competitor (Shah et al. 2024) that requires additional training. Our main contributions are summarized as follows:
  • We propose a novel framework called CoAR for AR customization. Unlike mainstream LoRA-based approaches, CoAR injects subject concepts into minimal context tokens without modifying pre-trained parameters.
@@ Introduction
  • We propose a training-free customization strategy that enables arbitrary subject–style combinations during inference. By concatenating subject and style context tokens, our method achieves high-quality stylization while preserving subject identity and contextual consistency without any additional training. Some generated samples are shown in Fig.˜1
  • Experimental results demonstrate that CoAR achieves competitive or superior performance on subject-driven personalized generation, while requiring substantially fewer trainable parameters. Furthermore, it achieves performance comparable to modern diffusion-based approaches in subject-style personalized generation.
Related Works
Autoregressive Image Generation
@@
In recent years, autoregressive (AR) image generation has rapidly evolved from pixel-level sequential modeling to token-level generation and further to multi-modal unified architectures. Early approaches (Van den Oord et al. 2016; Van Den Oord, Kalchbrenner, and Kavukcuoglu 2016) operated directly in the pixel space, which imposed heavy computational demands for generating high-resolution images. The VQ-VAE family (Ramesh et al. 2021; Van Den Oord, Vinyals et al. 2017) subsequently discretizes images into tokens, followed by training a Transformer (Vaswani et al. 2017) model using the Next Token Prediction (NTP) task to generate images. Recent research has explored two major directions: One is Visual autoregressive (VAR) models (Tian et al. 2024; Han et al. 2025). VAR models adopt a next-scale prediction framework that significantly improves image quality. The other is multi-modal unified AR models (Team 2024; Liu et al. 2024a; Xin et al. 2025). Chameleon (Team 2024) employs an early-fusion strategy to mix text and image tokens into a single sequence, enabling unified text-to-image, captioning, and inpainting generation. Lumina-mGPT (Liu et al. 2024a) further incorporates progressive FP-SFT, substantially enhancing high-resolution image synthesis. Lumina-mGPT 2.0 (Xin et al. 2025) approaches state-of-the-art diffusion models within a pure AR framework while maintaining strong multi-task generalization. These unified multi-modal AR models demonstrate strong potential as a foundation for versatile and scalable vision-language generation systems.
@@
Figure 2: Overview of the proposed CoAR framework for subject-driven personalization in multi-modal autoregressive models. (a) Layerwise Multimodal Context Learning, where learnable context tokens are injected into multiple Transformer layers for concept representation. (b) Dual Prior Preservation (DPP) regularizes the customized distribution against the pre-trained model to mitigate overfitting and language drift. (c) Context-Aware Self-Regularization (CASR) initializes and constrains context tokens towards the subject embedding space to enhance fidelity and re-contextualization. (d) Training-free subject–style composition by directly combining subject and style tokens to enable flexible customized generation.
Subject-Driven Image Generation
Subject-driven image generation aims to capture the concept of a specific subject from a limited set of reference images and generalize it to novel contexts. Existing approaches can broadly be categorized into diffusion-based and AR-based methods. Diffusion-based approaches (Ruiz et al. 2023; Chen et al. 2023a; Han et al. 2023; Kumari et al. 2023; Voynov et al. 2023; Gal et al. 2022) typically adapt the distribution of a pre-trained model to that of the target subject via either full fine-tuning or low-rank adaptation (LoRA) (Hu et al. 2022). During inference, a special identifier is employed to sample from the adapted distribution, allowing for customized generation. As the potential of AR models in image generation continues to be exploited, recent studies (Chung et al. 2025; Zheng et al. 2025; Sun et al. 2025; Wu et al. 2025) have begun to investigate subject-driven customization specifically tailored for AR models. For VAR-based models, ARBooth (Chung et al. 2025) introduces selective layer tuning to reduce complexity and employs prior distillation to mitigate the issue of language drift. FreeLoRA (Zheng et al. 2025) proposes a training-free multi-subject customization approach for VAR models. For multimodal AR models, PersonalAR (Sun et al. 2025) validates the combination of DreamBooth (Ruiz et al. 2023) and TI (Gal et al. 2022) within a LoRA framework on Lumina-mGPT. Proxy-Tuning (Wu et al. 2025) further enhances capability in subject-driven image generation by leveraging diffusion models as proxies.
These approaches follow the diffusion-based customization paradigm, adapting the generation backbone to specific subject distributions. A concurrent work, Yo’Chameleon (Nguyen et al. 2025), also introduces context optimization for multimodal generation of the auto-regressive model but suffers from overfitting and language drift, leading to reduced quality and subject fidelity. In contrast, CoAR represents subject or style information using a compact set of learnable context tokens. This eliminates the need to update backbone parameters while preserving the model’s generative capacity. By decoupling customization from backbone adaptation, CoAR offers an effective alternative to prior approaches.
Methodology
@@
In this work, we propose CoAR, a simple yet effective framework for subject-driven customization in multi-modal autoregressive (AR) models. The overall CoAR framework is shown in Fig. 2, consisting of:
@@
Preliminary
We first introduce the AR: in multimodal autoregressive (AR) text-to-image generation, both text and images are represented as sequences of discrete tokens drawn from a shared vocabulary $\mathcal{V}$. A text prompt is encoded as:
@@
where $\mathbf{x}_{<t}=\{\mathbf{x}_{1},\dots,\mathbf{x}_{t-1}\}$ and $\theta$ are the model parameters.
Training follows the next-token prediction paradigm, minimizing the negative log-likelihood of the ground-truth token sequence conditioned on the text prompt:
@@
Layerwise Multimodal Context Learning Strategy
We introduce a Layerwise Multimodal Contextual Learning (LMCL) strategy to capture the target subject concept while keeping the pre-trained model entirely frozen, requiring only a small number of learnable parameters for effective customization.
Specifically, for all images from the same subject, we define a shared set of multimodal learnable tokens as the representation of the subject concept:
@@
$\mathbf{U}_{i}^{\prime}=\{\mathbf{y}_{1},\mathbf{y}_{2},\dots,\mathbf{p}_{{[v]}}^{(i)},\dots,\mathbf{y}_{L},\mathbf{p}_{{I}}^{(i)},\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{T}\},$ (7)
where $\mathbf{p}_{{[v]}}^{(i)}$ replaces the placeholder [V] in the text sequence, and $\mathbf{p}_{{I}}^{(i)}$ is inserted before the first image token.
Finally, we optimize the representation of these multimodal tokens using Eq. (4) to progressively inject the concept of the specific subject into them. By learning only $\{\mathbf{p}_{[v]},\mathbf{p}_{I}\}$ while keeping the backbone frozen, LMCL achieves efficient subject-driven customization with minimal parameter overhead.
Dual Prior Preservation
To mitigate language drift and encourage generation diversity, we propose Dual Prior Preservation (DPP) to make the prior preservation introduced in DreamBooth (Ruiz et al. 2023) compatible with training of the multimodal AR model. We first generate a small set (6–8) of class images using the pre-trained model with the prompt “a photo of a [Class]”. These generated images, along with their corresponding text prompts, are tokenized into an input sequence $\mathbf{U}_{{cls}}$. We then perform a dual forward pass on $\mathbf{U}_{{cls}}$:
@@
Context-Aware Self-Regularization
@@
Method DINO CLIP-I CLIP-T Trainable Params Training Stage
Real Imges 0.774 0.885 N/A N/A N/A
TI (Ruiz et al. 2023) 0.569 0.780 0.255 N/A
DreamBooth(SD) (Ruiz et al. 2023) 0.668 0.803 0.305 N/A
DreamBooth (Imagen) (Ruiz et al. 2023) 0.696 0.812 0.306 N/A One-Stage
DreamBooth (Lumina-mGPT) (Wu et al. 2025) 0.5088 0.6752 0.2956 N/A
DreamBooth (Lumina-mGPT)+LoRA (Wu et al. 2025) 0.5338 0.6974 0.2652 N/A
PersonalAR (Lumina-mGPT)+LoRA (Sun et al. 2025) 0.657 0.778 0.312 50.4M
PersonalAR (Lumina-mGPT) (Sun et al. 2025) 0.671 0.805 0.302 1610.6M Two-Stage
Proxy-Tuning (Lumina-mGPT/SD3.5)+LoRA (Wu et al. 2025) 0.7523 0.8091 0.3123 142.6M
CoAR (Lumina-mGPT) (Ours) 0.7226 0.8151 0.3184 0.073M One-Stage
Table 1: Performance comparison of our CoAR and the recent competitors on the DreamBench dataset (Ruiz et al. 2023). The best results are marked in bold.
Figure 3: Qualitative comparison of subject-driven personalization on the DreamBench benchmark, selecting PersonalAR and Proxy-Tuning for comparison. The results for both competitors are taken directly from their original publications.
Customize Any Subject in Any Style

### v2

# DCoAR: Deep Concept Injection into Unified Autoregressive Models for Personalized Text-to-Image Generation

@@ Abstract
The unified autoregressive (AR) model excels at multimodal understanding and generation. However, its full potential in the domain of customized image generation has yet to be fully realized. Existing customization approaches for unified AR models face a fundamental dilemma: adaptation-based methods suffer from overfitting and scalability bottlenecks, while concept-injection paradigms are constrained by a shallow injection strategy that leads to poor visual fidelity and impaired re-contextualization. To address this, we propose DCoAR, a novel deep concept injection framework that maintains a completely frozen pre-trained model. DCoAR deeply integrates new concepts through a Layer-wise Multimodal Context Learning (LMCL) strategy, which is stabilized by a multi-faceted regularization scheme: a Dual Prior Preservation (DPP) loss to mitigate semantic drift and a Context-Aware Self-Regularization (CASR) loss to enhance re-contextualization. The framework also enables training-free subject customization in user-provided styles. Experiments demonstrate that DCoAR significantly outperforms previous injection-based methods and achieves performance competitive with adaptation-based approaches while requiring substantially fewer trainable parameters.
@@
Figure 1: Visualization of samples generated by our proposed DCoAR, showcasing subject personalization (left), where a few reference images are provided, and style personalization (right), where the subject is rendered in user-specified artistic styles without additional training.
1 Introduction
Unified multimodal autoregressive (AR) models [team2024chameleon, lumina_mgpt, mmr1, stylear, editar, acdit], have established a powerful paradigm for processing interleaved text and image data, demonstrating remarkable success across a range of generative and understanding tasks. Despite their success, the application of these models to personalized image generation has received comparatively less attention, leaving their full potential in this specific task open for exploration. Personalized image generation[chen2023subject, ruiz2023dreambooth, li2023blip, shin2025large, liu2024llm4genleveragingsemanticrepresentation, mosaicmultisubjectpersonalizedgeneration, TFcustom, she2025customvideox3dreferenceattention] aims to generate images of a specific subject or concept in new scenes, given only a few reference images. This capability has found applications in virtual try-on [ye2025slot, zhang2024two], personalized content creation [liu2024make], and advertising design [chen2023subject].
Personalization for unified AR models remains fundamentally challenging, and existing approaches fall short in ways that are increasingly incompatible with real-world demands. On the one hand, fine-tuning–based methods, inspired by DreamBooth [ruiz2023dreambooth, chen2023disenbooth, han2023svdiff, customdiff, OmniControl, EasyControl] and implemented through PEFT techniques such as LoRA [Prefixtuning, coop, hu2022lora, personalAR, wu2025proxytuning], attempt to rewrite model weights for each new concept. Yet AR models are far more brittle than diffusion models: adapting millions of parameters using only 3–5 images almost inevitably causes overfitting, catastrophic forgetting, and distortion of pre-trained priors. Worse still, maintaining a separate set of weights for every concept leads to unbounded storage costs, making this strategy fundamentally non-scalable for consumer-facing or large-scale personalization systems. On the other hand, token-injection approaches [UniCTokens, yoChameleon] promise scalability by keeping the backbone frozen, but they suffer from a deeper structural flaw. Injecting concept information only at the input layer forces the signal to pass through dozens of transformer blocks without reinforcement. This shallow injection bottleneck results in severe semantic attenuation, where fine-grained identity cues vanish as depth increases, and the model fails to bind the concept to complex prompts. Consequently, these methods struggle with visual fidelity, contextual adaptability, and semantic consistency, falling short of fine-tuning counterparts.
To overcome the constraints of shallow concept injection, we introduce Deep Concept Injection for unified AR models (DCoAR), an efficient and extensible framework for customized image generation. The core of our DCoAR framework is the Layer-wise Multimodal Context Learning (LMCL) strategy that encodes subject-specific information into a compact set of learnable tokens and injects them across multiple transformer layers, enabling deeper semantic alignment and high-fidelity feature propagation without modifying the frozen backbone. To ensure the robustness and quality of this deep injection process, DCoAR incorporates a multi-faceted regularization scheme consisting of two components. First, a Dual Prior Preservation (DPP) loss constrains the model against semantic drift via constraining the concept-conditioned features to stay close to the model’s native feature distribution while penalizing deviations from its latent feature gradients, ensuring that personalization does not compromise the inherent generative priors of the base AR model. Second, a Context-Aware Self-Regularization (CASR) enhances re-contextualization by preventing the concept tokens from over-specializing to the limited training prompts. CASR enforces cross-context consistency by aligning the model’s internal representations under both original and perturbed contexts, encouraging the tokens to stay flexible rather than memorizing appearance details. It further penalizes representation collapse across layers, ensuring that the injected concept signals remain discriminative and generalizable during deep propagation.
Our main contributions are summarized as follows:
  • We propose a deep concept injection framework for unified AR models. Unlike prior methods constrained by shallow, early-layer conditioning, DCoAR introduces a Layer-wise Multimodal Context Learning (LMCL) strategy that injects concept signals throughout the transformer stack, enabling persistent feature reinforcement and richer cross-modal interaction. This design achieves robust concept propagation and high-fidelity personalization without modifying a single parameter of the frozen backbone.
@@ 1 Introduction
  • We introduce a training-free subject–style composition strategy. By directly concatenating the learned context tokens for a subject and a style, DCoAR enables arbitrary, plug-and-play combinations at inference time, achieving high-quality, context-aware stylization without any additional optimization or fine-tuning. Some generated samples are shown in Fig.˜1.
  • Experimental results show that DCoAR sets a new state of the art for unified AR model personalization. Furthermore, DCoAR delivers strong gains in subject fidelity, instruction alignment, and parameter efficiency, requiring fewer than 0.1M trainable parameters. For stylization, our training-free pipeline produces high-fidelity results that rival modern LoRA-based methods without any additional training.
2 Related Works
2.1 Autoregressive Image Generation
@@
In recent years, autoregressive (AR) image generation has rapidly evolved from pixel-level sequential modeling to token-level generation and further to multi-modal unified architectures. Early approaches [pixelcnn, pixelrnn] operated directly in the pixel space, which imposed heavy computational demands for generating high-resolution images. The VQ-VAE family [dalle, vqvae] subsequently discretizes images into tokens, followed by training a Transformer [transformer] model using the Next Token Prediction (NTP) task to generate images. Recent research has explored two major directions: One is Visual autoregressive (VAR) models [var, infinity]. VAR models adopt a next-scale prediction framework that significantly improves image quality. The other is multi-modal unified AR models [team2024chameleon, lumina_mgpt, lumina2]. Chameleon [team2024chameleon] employs an early-fusion strategy to mix text and image tokens into a single sequence, enabling unified text-to-image, captioning, and inpainting generation. Lumina‑mGPT [lumina_mgpt] further incorporates progressive FP-SFT, substantially enhancing high-resolution image synthesis. Lumina‑mGPT 2.0 [lumina2] approaches state-of-the-art diffusion models within a pure AR framework while maintaining strong multi-task generalization. These unified multi-modal AR models demonstrate strong potential as a foundation for versatile and scalable vision-language generation systems.
@@
Figure 2: Overview of the proposed DCoAR framework for subject-driven personalization in multi-modal autoregressive models. (a) Layerwise Multimodal Context Learning, where learnable context tokens are injected into multiple Transformer layers for concept representation. (b) Dual Prior Preservation (DPP) regularizes the customized distribution against the pre-trained model to mitigate overfitting and language drift. (c) Context-Aware Self-Regularization (CASR) initializes and constrains context tokens towards the subject embedding space to enhance fidelity and re-contextualization. (d) Training-free subject–style composition by directly combining subject and style tokens to enable flexible customized generation.
2.2 Personalized Image Generation
Personalized image generation aims to capture a specific concept from a few reference images and render it in novel contexts. While this task has been extensively studied in diffusion models [ruiz2023dreambooth, chen2023disenbooth, han2023svdiff, customdiff, pplus_TI, textualinversion], its application to unified autoregressive (AR) models presents unique challenges and has led to the emergence of two primary paradigms. The first is the adaptation-based paradigm, which employs parameter-efficient fine-tuning (PEFT) like LoRA [hu2022lora] to adapt a model’s weights for a new concept [personalAR, wu2025proxytuning, ARBooth, freelora]. While capable of high fidelity, this approach is hampered by overfitting and scalability bottlenecks for large-scale deployment. The second is the concept-injection paradigm, which inserts concepts into a frozen model via learnable tokens to ensure scalability. Pioneering works in this area, such as Yo’Chameleon [yoChameleon] and UniCTokens [UniCTokens], have explored different token designs. However, these methods are fundamentally constrained by a shallow injection strategy, where confining the concept to the input layer limits the propagation of visual details, resulting in a fidelity gap compared to adaptation-based approaches. In contrast, DCoAR advances this paradigm by introducing a deep injection strategy. Through Layer-wise Multimodal Context Learning (LMCL), concept tokens are injected across multiple transformer layers, ensuring robust feature propagation. This approach effectively bridges the fidelity gap with adaptation-based methods while preserving the scalability inherent to the concept-injection paradigm.
3 Method
@@
In this work, we propose DCoAR, a simple yet effective framework for subject-driven customization in unified autoregressive (AR) models. The overall DCoAR framework is shown in Fig. 2, consisting of:
@@
3.1 Preliminary
In unified autoregressive (AR) text-to-image generation, both text and images are represented as sequences of discrete tokens drawn from a shared vocabulary $\mathcal{V}$. A text prompt is encoded as:
@@ 3.1 Preliminary
where $\mathbf{x}_{<t}=\{\mathbf{x}_{1},\dots,\mathbf{x}_{t-1}\}$ and $\theta$ are the model parameters. Training follows the next-token prediction paradigm, minimizing the negative log-likelihood of the ground-truth token sequence conditioned on the text prompt:
@@
3.2 Layerwise Multimodal Context Learning Strategy
We introduce a Layerwise Multimodal Contextual Learning (LMCL) strategy to capture the target subject concept while keeping the pre-trained model entirely frozen, requiring only a small number of learnable parameters for effective customization. Specifically, for all images from the same subject, we define a shared set of multimodal learnable tokens as the representation of the subject concept:
@@ 3.2 Layerwise Multimodal Context Learning Strategy
$\mathbf{U}_{i}^{\prime}=\{\mathbf{y}_{1},\mathbf{y}_{2},\dots,\mathbf{p}_{[v]}^{(i)},\dots,\mathbf{y}_{L},\mathbf{p}_{I}^{(i)},\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{T}\},$
where $\mathbf{p}_{{[v]}}^{(i)}$ replaces the placeholder [V] in the text sequence, and $\mathbf{p}_{{I}}^{(i)}$ is inserted before the first image token. Finally, we optimize the representation of these multimodal tokens using Eq. (4) to progressively inject the concept of the specific subject into them.
By learning only $\{\mathbf{p}_{[v]},\mathbf{p}_{I}\}$ while keeping the backbone frozen, LMCL achieves efficient subject-driven customization with minimal parameter overhead.
3.3 Dual Prior Preservation
To mitigate language drift and encourage generation diversity, we propose Dual Prior Preservation (DPP) to make the prior preservation introduced in DreamBooth [ruiz2023dreambooth] compatible with training of the multimodal AR model. We first generate a small set (6–8) of class images using the pre-trained model with the prompt “a photo of a [Class]”. These generated images, along with their corresponding text prompts, are tokenized into an input sequence $\mathbf{U}_{{cls}}$. We then perform a dual forward pass on $\mathbf{U}_{{cls}}$:
@@ 3.3 Dual Prior Preservation
3.4 Context-Aware Self-Regularization
@@ 3.5 Overall Objective
Figure 3: Qualitative comparison of subject-driven personalization on the DreamBench benchmark. As concept-injection methods generally exhibit lower visual fidelity, we focus the qualitative comparison on the more competitive adaptation-based methods. Additional comparisons are available in the supplementary material.
Method Tyeps Method DINO CLIP-I CLIP-T Trainable Params
Real Imges 0.774 0.885 N/A N/A
Diffusion Models TI[textualinversion] 0.569 0.780 0.255 N/A
DreamBooth(SD)[ruiz2023dreambooth] 0.668 0.803 0.305 N/A
DreamBooth (Imagen)[ruiz2023dreambooth] 0.696 0.812 0.306 N/A
DreamBooth (Flux)[wu2025proxytuning] 0.732 0.802 0.314 N/A
Unified AR Models DreamBooth(Lumina-mGPT)[wu2025proxytuning] 0.509 0.675 0.295 N/A
DreamBooth(Lumina-mGPT)[wu2025proxytuning] +LoRA 0.534 0.697 0.265 N/A
Yo’Chameleon[yoChameleon] 0.542 0.795 0.225 0.13M
UniCTokens[UniCTokens] 0.599 0.782 0.304 0.13M
PersonalAR[personalAR] 0.671 0.805 0.302 1610.6M
PersonalAR[personalAR] +LoRA 0.657 0.778 0.312 50.4M
Proxy-Tuning[wu2025proxytuning] 0.752 0.809 0.312 142.6M
DCoAR(Ours) 0.723 0.815 0.318 0.073M
Table 1: Performance comparison with state-of-the-art methods on the DreamBench dataset. The results for competing methods are directly cited from previous publications and details on parameter count calculation can be found in the supplementary material.
3.6 Customize Any Subject in Any Style